# Positive microlocal holonomies are globally regular

### v1

@@ 3.1.2. Two properties of the Kashiwara-Schapira stack
We need the following two results on the $\infty$-sheaf $\operatorname{\operatorname{\mu Sh}}$: Theorem 3.11 and Proposition 3.13. The former states invariance of the stack under non-characteristic deformations of the Lagrangian subsets, see e.g.  [Nad15, NS20, Zho23]. Such deformations include Liouville homotopies of Weinstein manifolds or pairs, see e.g.  [NS20, Thm. 9.14]. (Confer also [Li, Thm. 1.1 & Rem. 1.5] or, alternatively, it follows from [GPS24].)
Theorem 3.11 ([NS20, Thm. 9.14] or [Li, Thm. 1.1]).
@@
$\operatorname{Sh}_{\Lambda_{f}}(\mathbb{R}^{2})_{0}\cong\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}_{\tilde{f}}}(\mathbb{L}_{\tilde{f}}),$
where the subscript 00 denotes the full subcategory of sheaves with bounded supports in $\mathbb{R}^{2}$. Different choices of real Morsifications lead to different Lagrangian skeleta. Theorem 3.11 implies that the global sections of the Kashiwara-Schapira stack on these skeleta are all equivalent: they are isomorphic to $\operatorname{Sh}_{\Lambda_{f}}(\mathbb{R}^{2})_{0}$.
$(ii)$ More generally, let $\Lambda_{\beta}\subset(\mathbb{R}^{3},\xi_{\text{st}})$ be the $(-1)$-closure of a positive braid $\beta$, cf. [CN22, Section 2]. Suppose that the Lusztig cycles $\mathscr{D}_{\mathfrak{w}}$ of a weave $\mathfrak{w}$ for $\beta\delta(\beta)$ provide a complete $\mathbb{L}$-compressing system with Lagrangian disks $\mathscr{D}^{*}_{\mathfrak{w}}$, e.g. if $\beta$ contains a $w_{0}$ subword, and let $L_{\mathfrak{w}}$ be the filling of $\Lambda_{\beta}$ associated to $\mathfrak{w}$. See [CZ22, Section 2], [CW24, Sections 2&3] and [CGG+22, Section 4] for details on weaves and $\mathbb{L}$-compressing systems. Then the union $\mathbb{L}_{\mathfrak{w}}:=L_{\mathfrak{w}}\cup\mathscr{D}_{\mathfrak{w}}^{*}$ is a relative arboreal Lagrangian skeleton for $(\mathbb{R}^{4},\Lambda)$. In this case
@@
Note that there are many choices for $\mathfrak{w}$ given a $\beta$. Typically infinitely many if one considers $\beta$ cyclically. In either case, there are many choices of Lagrangian skeleta $\mathbb{L}_{\mathfrak{w}}$ for a fixed such $\Lambda_{\beta}$. By Theorem 3.11, the global sections of their Kashiwara-Schapira stacks are all equivalent: they are isomorphic to $\operatorname{Sh}_{\Lambda_{\beta}}(\mathbb{R}^{2})_{0}$, which is independent of $\mathfrak{w}$. $\Box$
@@ 3.2. Corestriction functors for Kashiwara-Schapira stacks
By genericity and invariance, it suffices to consider the case where the Legendrian embedding $\mathbb{L}\subset T^{*}_{\infty}M$ is such that the front projection $\pi:\mathbb{L}\longrightarrow M$ is finite-to-one. We claim that there exists a topological basis of $T^{*}_{\infty}M$ on which $\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}}$ takes values in $\mathrm{Pr}^{\mathrm{R}}_{\omega,\mathrm{dg},k}$, the $\infty$-category whose objects are compactly generated dg-categories and whose functors preserve limits and (filtered) colimits, cf.  [Lur09, Def. 5.5.7.5]. The claim is proven as follows. By Remark 3.6, when $\pi:\mathbb{L}\to M$ is finite-to-one, for any point in $T^{*}M$, there exists an open ball $B\subset M$ together with a union of open balls $\Omega\subset T^{*}_{\infty}B$ such that $\mathbb{L}\cap T^{*}_{\infty}B\subset\Omega$, and
@@
Since the restriction functor on all sheaves $r^{*}:\operatorname{Sh}(B)\longrightarrow\operatorname{Sh}(B^{\prime})$ preserves limits and colimits (whose left adjoint is $r_{!}$ and right adjoint is $r_{*}$), on the full subcategories $r^{*}:\operatorname{Sh}_{\mathbb{L}}(B)\longrightarrow\operatorname{Sh}_{\mathbb{L}}(B^{\prime})$ it also preserves limits and colimits and thus admits left and right adjoints. (See [Kuo23, Thm. 1.2] for explicit formulae for these adjoints.) Since the restriction also preserves local systems, the adjunction descends to localizations $r^{*}:\operatorname{Sh}_{\mathbb{L}}(B)/\mathrm{Loc}(B)\to\operatorname{Sh}_{\mathbb{L}}(B^{\prime})/\mathrm{Loc}(B^{\prime})$. This shows that, for this basis, $\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}}$ sends morphisms to those in $\mathrm{Pr}^{\mathrm{R}}_{\omega,\mathrm{dg},k}$ and thus establishes the claim.
@@
Via the equivalence between microlocal sheaves and partially wrapped Fukaya categories, established in [GPS24], we expect that in good situations, the corestriction functor for open inclusions of stratified Lagrangian subsets coincides with the pushfoward functor for proper inclusions of Weinstein sectors, cf. [GPS20]. In line with [Kuo23], we also expect that corestriction functors can be characterized by sheaf-theoretic wrappings. $\Box$
@@ 3.3. Corestriction functors in the pointed setting
${\displaystyle\prod_{i=1}^{|\mathfrak{t}|}\operatorname{\mbox{Mod}}(k)}$${\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}}(\mathbb{L})}$${\operatorname{\mbox{Mod}}(k)}$$\scriptstyle{m_{\mathbb{L},\mathfrak{t}}^{\ell}}$$\scriptstyle{\Delta^{\ell}}$
@@
An alternative to Corollary 3.17 is to give a topological model for the category of sheaves with decorations $\operatorname{Sh}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$ via Lagrangian skeleta. The construction reads as follows:
Proposition 3.18.
Let $\Lambda\subset(T^{*}_{\infty}\mathbb{R}^{2},\xi_{st})$ be a Legendrian link and $\mathfrak{t}\subset\Lambda$ be a set of basepoints, with at least one basepoint per component, and $\mathbb{L}$ a relative Lagrangian skeleton of the Weinstein pair $(T^{*}\mathbb{R}^{2},\Lambda)$. Consider a Lagrangian skeleton $\mathbb{L}_{\mathfrak{t}}$ obtained by gluing (the boundary of) one Lagrangian disk $\mathbb{D}$ to $\mathbb{L}$ along the $|\mathfrak{t}|$ basepoints, using small neighbourhoods of the basepoints to obtain an arboreal Lagrangian skeleton. Then there is an equivalence
$\operatorname{Sh}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}\cong\operatorname{\operatorname{\mu Sh}}_{\mathbb{L},\mathfrak{t}}(\mathbb{L},\mathfrak{t})\cong\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}_{\mathfrak{t}}}(\mathbb{L}_{\mathfrak{t}}).$
Proof.
Let $i:\mathfrak{t}\hookrightarrow\mathbb{D}$ be the embedding of the basepoints in the boundary $\partial\mathbb{D}$. Since $\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}_{\mathfrak{t}}}$ is a sheaf and cosheaf of dg-categories under the restriction and corestriction functors, after using Theorem 3.11 and Proposition 3.13, we have a homotopy colimit diagram
${\mathrm{Loc}(\mathfrak{t})}$${\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}}$${\mathrm{Loc}(\mathbb{D})}$${\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}_{\mathfrak{t}}}(\mathbb{L}_{\mathfrak{t}}).}$$\scriptstyle{m_{\Lambda,\mathfrak{t}}^{\ell}}$$\scriptstyle{i^{*\ell}}$
Consider the equivalence $\mathrm{Loc}(\mathbb{D})\cong\operatorname{Mod}(k)$ and $\mathrm{Loc}(\mathfrak{t})\cong\prod_{i=1}^{|\mathfrak{t}|}\operatorname{Mod}(k)$. Then the restriction functor $i^{*}:\mathrm{Loc}(\mathbb{D})\longrightarrow\mathrm{Loc}(\mathfrak{t})$ agrees with the diagonal map
$\Delta:\operatorname{Mod}(k)\longrightarrow\displaystyle\prod_{i=1}^{|\mathfrak{t}|}\operatorname{Mod}(k).$
Therefore, the corestriction functor $i^{*\ell}:\mathrm{Loc}(\mathfrak{t})\longrightarrow\mathrm{Loc}(\mathbb{D})$ agrees with the left adjoint
$\Delta^{\ell}:\displaystyle\prod_{i=1}^{|\mathfrak{t}|}\operatorname{Mod}(k)\longrightarrow\operatorname{Mod}(k)$
by taking coproducts. Since $\operatorname{Sh}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$ is the homotopy colimit of the diagram
${\prod\nolimits_{i=1}^{|\mathfrak{t}|}\operatorname{Mod}(k)}$${\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}}$${\operatorname{Mod}(k)}$${\operatorname{Sh}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}}$$\scriptstyle{m_{\Lambda,\mathfrak{t}}^{\ell}}$$\scriptstyle{\Delta^{\ell}}$
this proves the required equivalence $\operatorname{Sh}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}\cong\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}_{\mathfrak{t}}}(\mathbb{L}_{\mathfrak{t}})$. ∎
Remark 3.19.
In Proposition 3.18 we did not specify how $\mathbb{L}$ and $\mathbb{D}$ are glued along neighbourhoods of the basepoints: specifically, we did not specify an orientation of $\Lambda$ and a cyclic order of the basepoints $\mathfrak{t}\subset\Lambda$. However, as is illustrated in the proof, the resulting category $\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}_{\mathfrak{t}}}(\mathbb{L}_{\mathfrak{t}})$ does not depend on the way $\mathbb{L}$ and $\mathbb{D}$ are glued together and is isomorphic to $\operatorname{Sh}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$. That said, in practice, when $\mathbb{L}$ is the Lagrangian skeleton of an $\mathbb{L}$-compressing system for an oriented Lagrangian filling, we always make the technical assumption that $\mathbb{L}$ and $\mathbb{D}$ are glued by an orientation-preserving map on the neighbourhood of the basepoints in order to make sure the relative cycles extend to absolute cycles that are co-oriented (so that we can readily use Proposition 3.13). $\Box$
@@ 3.4. Lagrangians fillings and Kashiwara-Schapira stacks
which preserves limits and colimits and thus admits left and right adjoints. Together with the compact generation of the categories, cf. [GPS24, Cor. 4.22], this shows that on the topological basis the inclusion functor lands in the $\infty$-category $\operatorname{Pr}^{\text{R}}_{\omega,dg,k}$. Since $\operatorname{Pr}^{\text{R}}_{\omega,dg,k}$ admits all small limits, for global sections, the inclusion functor also lands in $\operatorname{Pr}^{\text{R}}_{\omega,dg,k}$ and the statement follows. ∎
@@
The argument in [GPS24, Thm. 4.13] shows that the localization functor $\iota^{*}:\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}}(\mathbb{L})\hookrightarrow\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}^{\prime}}(\mathbb{L}^{\prime})$ is given by localization along all the corepresentatives of microstalk functors on $\mathbb{L}\setminus\mathbb{L}^{\prime}$. From the perspective of wrapped Fukaya categories, the localization functor can be viewed as a Viterbo restriction functor from $X$ to a Weinstein neighbourhood of the Lagrangian $L$ (symplectomorphic to a disk cotangent bundle $T^{*}_{<\varepsilon}L$, $\varepsilon\in\mathbb{R}_{+}$ small enough). $\Box$
@@
It can be proven that the inclusion functor $\iota_{*}:\mathrm{Loc}(L)\hookrightarrow\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}}(\mathbb{L})$ agrees with the sheaf quantization functor $\mathrm{Loc}(L)\hookrightarrow\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$ from [NS20] under the equivalence $\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}}(\mathbb{L})\cong\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$, cf. [Li, Thm. 1.1]. One can also show that the sheaf quantization functors of [NS20] and [JT24] agree in simple cases, cf. [Li23, Prop. 1.8] and [CL23, Rem. B.23]. The latter functor is the sheaf quantization functor used to define algebraic torus charts and microlocal holonomies in the moduli of sheaves $\mathfrak{M}(\Lambda)$ and $\mathfrak{M}(\Lambda,\mathfrak{t})$, cf. [CW24, Section 4.3] and [CL23, Section 5.1]. $\Box$
@@
${\displaystyle\prod_{i=1}^{|\mathfrak{t}|}\operatorname{\mbox{Mod}}(k)}$${\mathrm{Loc}(L)}$${\operatorname{\mbox{Mod}}(k)}$$\scriptstyle{(i_{\mathfrak{t}}^{*})^{\ell}}$$\scriptstyle{\Delta^{\ell}}$
@@ 4.1.1. Some context
As it is apparent from our previous works, see e.g.  [CGG+22, CW24], it can be challenging to directly construct such regular functions on $\mathfrak{M}({\Lambda,\mathfrak{t}})$. Thus far, the only successful strategy has been as follows:
@@
    Employ ad-hoc methods, often exploiting the specific combinatorics (such as plabic graphs in [CW24] or weaves in [CGG+22]) and relation to cluster mutations, so as to argue that the regular functions built on $T(L)$ extend to regular functions on $\mathfrak{M}({\Lambda,\mathfrak{t}})$. At core, this is an argument that requires explicit formulas related to the underlying combinatorics. See for instance [CW24, Section 4.9 & Prop. 4.38] or [CGG+22, Section 5.3], specifically Lemmas 5.26, 5.27 and 5.28 in loc. cit.
@@
Remark 4.1.
The principle of conservation of difficulty is not violated: the work is instead found in the results on existence and properties of certain categories and functors between them and the few key computations in this article. Overall, these include the results developed within the framework of homotopical algebraic geometric, cf.  [TV05, TV08, TV07], the existence of corestriction functors and the properties of the Kashiwara-Schapira stack in microlocal sheaf theory, cf.  [Gui23, GKS12, KS90, Nad16], the computations in Section 3 and the results in the subsections below. $\Box$

### v2

@@ 3.1.2. Two properties of the Kashiwara-Schapira stack
We need the following two results on the $\infty$-sheaf $\operatorname{\operatorname{\mu Sh}}$: Theorem 3.13 and Proposition 3.15. The former states invariance of the stack under non-characteristic deformations of the Lagrangian subsets, see e.g.  [NAD15, NS20, ZHO23]. Such deformations include Liouville homotopies of Weinstein manifolds or pairs, see e.g.  [NS20, Thm. 9.14]. (Confer also [55, Thm. 1.1 & Rem. 1.5] or, alternatively, it follows from [GPS24].)
Theorem 3.13 ([NS20, Thm. 9.14] or [55, Thm. 1.1]).
@@
[…]
where the subscript 0 denotes the full subcategory of sheaves with bounded supports in $\mathbb{R}^{2}$. Different choices of real Morsifications lead to different Lagrangian skeleta. Theorem 3.13 implies that the global sections of the Kashiwara-Schapira stack on these skeleta are all equivalent: they are quasi-equivalent to $\operatorname{Sh}_{\Lambda_{f}}(\mathbb{R}^{2})_{0}$.
$(ii)$ More generally, let $\Lambda_{\beta}\subset(\mathbb{R}^{3},\xi_{\text{st}})$ be the (1)(-1)-closure of a positive braid $\beta$, cf. [CN22, Section 2]. Suppose that the Lusztig cycles $\mathscr{D}_{\mathfrak{w}}$ of a weave $\mathfrak{w}$ for $\beta\delta(\beta)$ provide a complete $\mathbb{L}$-compressing system with Lagrangian disks $\mathscr{D}^{*}_{\mathfrak{w}}$, e.g. if $\beta$ contains a $w_{0}$ subword, and let $L_{\mathfrak{w}}$ be the filling of $\Lambda_{\beta}$ associated to $\mathfrak{w}$. See [CZ22, Section 2], [CW24, Sections 2&3] and [CGG+25, Section 4] for details on weaves and $\mathbb{L}$-compressing systems. Then the union $\mathbb{L}_{\mathfrak{w}}:=L_{\mathfrak{w}}\cup\mathscr{D}_{\mathfrak{w}}^{*}$ is a relative arboreal Lagrangian skeleton for $(\mathbb{R}^{4},\Lambda)$. In this case
@@
Note that there are many choices for $\mathfrak{w}$ given a $\beta$. Typically infinitely many if one considers $\beta$ cyclically. In either case, there are many choices of Lagrangian skeleta $\mathbb{L}_{\mathfrak{w}}$ for a fixed such $\Lambda_{\beta}$. By Theorem 3.13, the global sections of their Kashiwara-Schapira stacks are all equivalent: they are quasi-equivalent to $\operatorname{Sh}_{\Lambda_{\beta}}(\mathbb{R}^{2})_{0}$, which is independent of $\mathfrak{w}$. $\Box$
@@ 3.2. Corestriction functors for Kashiwara-Schapira stacks
By genericity and invariance, it suffices to consider the case where the Legendrian embedding $\mathbb{L}\subset T^{*}_{\infty}M$ is such that the front projection $\pi:\mathbb{L}\longrightarrow M$ is finite-to-one. We claim that there exists a topological basis of $T^{*}_{\infty}M$ on which $\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}}$ takes values in $\mathrm{Pr}^{\mathrm{R}}_{\omega,\mathrm{dg},k}$, the $\infty$-category whose objects are compactly generated dg-categories and whose functors preserve limits and (filtered) colimits, cf.  [LUR09, Def. 5.5.7.5]. The claim is proven as follows. By Lemma 3.7, when $\pi:\mathbb{L}\to M$ is finite-to-one, for any point in $T^{*}M$, there exists an open ball $B\subset M$ together with a union of open balls $\Omega\subset T^{*}_{\infty}B$ such that $\mathbb{L}\cap T^{*}_{\infty}B\subset\Omega$, and
@@
Since the restriction functor on all sheaves $r^{*}:\operatorname{Sh}(B)\longrightarrow\operatorname{Sh}(B^{\prime})$ admits left adjoint $r_{!}$ and right adjoint $r_{*}$, and the inclusion $\operatorname{Sh}_{\mathbb{L}}(B)\hookrightarrow\operatorname{Sh}(B)$ admits left and right adjoints by [KUO23, Thm. 1.2], we know that on the full subcategories the restriction $r^{*}:\operatorname{Sh}_{\mathbb{L}}(B)\longrightarrow\operatorname{Sh}_{\mathbb{L}}(B^{\prime})$ also admits left and right adjoints. Since the restriction also preserves local systems, the adjunction descends to localizations $r^{*}:\operatorname{Sh}_{\mathbb{L}}(B)/\mathrm{Loc}(B)\to\operatorname{Sh}_{\mathbb{L}}(B^{\prime})/\mathrm{Loc}(B^{\prime})$. This shows that, for this basis, $\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}}$ sends morphisms to those in $\mathrm{Pr}^{\mathrm{R}}_{\omega,\mathrm{dg},k}$ and thus establishes the claim.
@@
Via the equivalence between microlocal sheaves and partially wrapped Fukaya categories, established in [GPS24], we expect that in good situations, the corestriction functor for open inclusions of stratified Lagrangian subsets coincides with the pushfoward functor for proper inclusions of Weinstein sectors, cf. [GPS20]. In line with [KUO23, Theorem 1.2 & Section 3.2] and [GPS20, Section 3.5], we also expect that corestriction functors can be characterized by sheaf-theoretic wrappings, i.e. we expect that applying each such corestriction functor can be expressed as a certain colimit of a diagram of sheaves indexed by positive Hamiltonian isotopies of the Weinstein sector. This would be the symplectic topological incarnation of the pointwise colimit formula for left Kan extensions, see e.g. [LUR09, Prop. 4.3.3.10]. $\Box$
@@ 3.3. Corestriction functors in the pointed setting
$\prod_{i=1}^{|\mathfrak{t}|}\operatorname{\mbox{Mod}}(k)$$\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}}(\mathbb{L})$$\operatorname{\mbox{Mod}}(k)$$\scriptstyle{m_{\mathbb{L},\mathfrak{t}}^{\ell}}$$\scriptstyle{\Delta^{\ell}}$
@@ 3.4. Lagrangians fillings and Kashiwara-Schapira stacks
which admits left and right adjoints by [GPS24, Cor. 4.22]. Together with the compact generation of the categories, cf. [GPS24, Cor. 4.21], this shows that on the topological basis the inclusion functor lands in the $\infty$-category $\operatorname{Pr}^{\text{R}}_{\omega,dg,k}$. Since $\operatorname{Pr}^{\text{R}}_{\omega,dg,k}$ admits all small limits, for global sections, the inclusion functor also lands in $\operatorname{Pr}^{\text{R}}_{\omega,dg,k}$ and the statement follows. ∎
@@
The argument in [GPS24, Thm. 4.13 and Cor. 4.21] shows that the localization functor $\iota^{*}:\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}}(\mathbb{L})\hookrightarrow\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}^{\prime}}(\mathbb{L}^{\prime})$ is given by localization along all the corepresentatives of microstalk functors on $\mathbb{L}\setminus\mathbb{L}^{\prime}$. From the perspective of wrapped Fukaya categories, the localization functor can be viewed as a Viterbo restriction functor from $X$ to a Weinstein neighbourhood of the Lagrangian $L$ (symplectomorphic to a disk cotangent bundle $T^{*}_{<\varepsilon}L$, $\varepsilon\in\mathbb{R}_{+}$ small enough). $\Box$
@@
It can be proven that the inclusion functor $\iota_{*}:\mathrm{Loc}(L)\hookrightarrow\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}}(\mathbb{L})$ agrees with the sheaf quantization functor $\mathrm{Loc}(L)\hookrightarrow\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$ from [NS20] under the equivalence $\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}}(\mathbb{L})\cong\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$, cf. [55, Thm. 1.1]. One can also show that the sheaf quantization functors of [NS20] and [JT24] agree in simple cases, cf. [LI23, Prop. 1.8] and [CL23, Rem. B.23]. The latter functor is the sheaf quantization functor used to define algebraic torus charts and microlocal holonomies in the moduli of sheaves $\mathfrak{M}(\Lambda)$ and $\mathfrak{M}(\Lambda,\mathfrak{t})$, cf. [CW24, Section 4.3] and [CL23, Section 5.1]. $\Box$
@@
$\prod_{i=1}^{|\mathfrak{t}|}\operatorname{\mbox{Mod}}(k)$$\mathrm{Loc}(L)$$\operatorname{\mbox{Mod}}(k)$$\scriptstyle{(i_{\mathfrak{t}}^{*})^{\ell}}$$\scriptstyle{\Delta^{\ell}}$
@@ 4.1.1. Some context
As it is apparent from our previous works, see e.g.  [CGG+25, CW24], it can be challenging to directly construct such regular functions on $\mathfrak{M}({\Lambda,\mathfrak{t}})$. Thus far, the only successful strategy has been as follows:
@@
    Employ ad-hoc methods, often exploiting the specific combinatorics (such as plabic graphs in [CW24] or weaves in [CGG+25]) and relation to cluster mutations, so as to argue that the regular functions built on $T(L)$ extend to regular functions on $\mathfrak{M}({\Lambda,\mathfrak{t}})$. At core, this is an argument that requires explicit formulas related to the underlying combinatorics. See for instance [CW24, Section 4.9 & Prop. 4.38] or [CGG+25, Section 5.3], specifically Lemmas 5.26, 5.27 and 5.28 in loc. cit.